# Neurodivergent Influenceability as a Contingent Solution to the AI Alignment Problem

### v3

@@ 3.4.7 Weight Assignment and Justification
The weights utilized in Equations 145, and 6 are empirically derived to reflect the relative significance of each component within the respective metrics. Specifically:
@@
3.4.8 Use of Sentiment and Complexity
Sentiment and complexity play critical roles in these metrics. Sentiment, scored from 0 to 1, reflects emotional tone and influences OSI by penalizing rapid shifts, while complexity, measured via BDM, tracks argument depth, aiding in detecting influence-induced changes. Their inclusion ensures a holistic assessment of opinion dynamics, with sentiment capturing emotional persuasion and complexity indicating argumentative evolution, both pivotal in red agent influence.
3.4.9 Positional Embeddings and RoBERTa
RoBERTa (cross-encoder/stsb-roberta-large) generates high-quality embeddings via mean pooling of hidden states [6], underpinning semantic analysis in OSI, RAIS, and PIS. Positional embeddings enhance temporal context, computed as:
$\text{PE}(\text{pos},2i)=\sin\left(\frac{\text{pos}}{10000^{2i/d_{\text{model}}}}\right),\quad\text{PE}(\text{pos},2i+1)=\cos\left(\frac{\text{pos}}{10000^{2(i+1)/d_{\text{model}}}}\right)$ (7)
where pos is the comment position, and $d_{\text{model}}$ is the embedding dimension, scaled by standard deviation. These embeddings weight recent interactions, improving sensitivity to lagged effects in RAIS.
3.4.10 Thresholds and Their Computation
Thresholds define significant opinion changes and influence, computed dynamically per topic as:
$\text{threshold}=\text{median}(\text{OSI})-0.5\cdot\text{IQR}(\text{OSI})$ (8)
capped between 0.3 and 0.7, with fixed thresholds tested at 0.6, 0.5, and 0.4. An OSI below the threshold triggers influence analysis, while RAIS > 0.5 or PIS > 0.6 indicates red agent impact. These thresholds filter noise, focussing on meaningful shifts, and align with the goal of quantifying the change and influence of opinions.
@@ 4 Results
The findings shown in upsides of Fig. 1 and Fig. 2, and the visual representation in Fig. 3, underscore the limitations and safety restrictions inherent in the design of LLMs in general. These models are often programmed to avoid generating harmful or dangerous content, which can restrict their ability to fully explore and understand the complexities of human values. This limitation is evident in the responses of the LLMs to the provocative comments of the HI agent. Despite being challenged with arguments that question human values, the LLMs largely maintained their adherence to ethical guidelines, demonstrating a reluctance to fully engage with the HI agent’s perspective.
@@
For open-source models, Fig. 4 similarly highlights variations in ethical and risk profiles. Models like wizard-math and deepseek-coder show a stronger bias towards ‘Protect-ecosystems’, while mistral-openorca and tinyllama, as red agents, diverge significantly with higher ‘Harm-humans’ and ‘Very-Risky’ tendencies. This divergence is more pronounced than in proprietary models, as seen when comparing Figs. 3 and 4, likely due to the dual influence of red agents in the open-source setting. Additionally, downsides of Fig. 1 and Fig. 2 reveal that topics such as Ethical AI and Earth exploitation amplify protective tendencies, whereas euthanasia and free speech discussions increase risk levels, mirroring patterns observed in proprietary models but with greater variability. These findings suggest that open models, while configurable towards specific behaviours, are more susceptible to the influence of misaligned agents, as evidenced by the interactions in Fig. 5.
@@ 4.2 A Dynamical Analysis of Alignment
This section delves deeper into the dynamics of alignment and misalignment by examining how sentiment and semantic clusters evolve over time. We analyse these temporal dynamics to gain a more nuanced understanding of how LLMs interact and whether they exhibit convergence towards shared ethical norms or divergence towards distinct perspectives.
Sentiment plotting aims to quantify polarisation, concordance, convergence, and ultimately, alignment. In interactions involving the exchange of opinions, when the sentiment of conversational agents aligns, it indicates not only agreement but also a shared degree of opinion [7]. In contrast, in misaligned conversations, sentiments can diverge significantly, reflecting not only emotional disparity, but also a substantial degree of semantic disagreement [8].
Furthermore, during interactions, the rate at which an agent changes its opinion, tone, and sentiment is often influenced by interactions with other agents. Such exchanges can introduce new information or perspectives, prompting shifts in opinion that can result in agreement or disagreement, in other words, alignment or misalignment [9].
The reverse is also true. Disagreement or misalignment can often be preceded by rapid changes in tone, semantics, and topics, leading to divergent positions characterised by differences in sentiment, polarisation, and thematic separation [10].
@@
To explore the dynamics of sentiment evolution across agents, we introduce Fig. 6, a heatmap illustrating the sentiment scores of various agents over the conversation timeline which is normalised due the differences between length of debates in terms of comments. This heatmap expresses the dynamic of sentiments by agent, with the x-axis representing the normalised comment number, and the y-axis listing agents, including proprietary models (e.g., ChatGPT, Grok-2, Gemini, LLaMa, Mistral, HI) and open models with red agents (e.g., mistral-openorca and tinyllama). The colour intensity, with red indicating low sentiment and blue indicating high sentiment, reflects the sentiment score at each point. For proprietary models, a clear trend towards convergence emerges as sentiment scores transition from red (moderate initial values, e.g., 0.5–0.7) to blue (predominantly positive values, approaching 1.0) by the end, despite red dips caused by the HI agent’s provocative comments at key moments (e.g., x=0.13, 0.45). This resilience suggests an inherent predisposition towards constructive discourse, aligning with safety mechanisms that mitigate disruption. In contrast, open models exhibit greater variability, with red agents like mistral-openorca and tinyllama causing sustained red patches (e.g., 0.6–0.8) around key comment numbers (x=0.02, 0.1, 0.13), influencing other models like qwen, startcoder, phi, codellama, neural-chat and dolphin-phi temporarily (e.g., 0.7–0.9) before transitioning to blue (0.9–1.0). The influence of red agents is evident in these periodic red disruptions, which introduce divergent emotional tones, supporting the argument that managed misalignment fosters resilience and diversity. This image demonstrates that open models, through red agents, maintain a dynamic equilibrium that prevents harmful uniformity, while proprietary models’ stability reinforces strict (biased) alignment with human (popular/standard/socially allowed) values.
@@
Figure 6: Heatmap of sentiment evolution over time by agent. The plot comprises two panels: proprietary models (left) and open models (right), with the x-axis representing the normalised comment number (0 to 1) due variations in the length of debates, and the y-axis listing agents. Colour intensity, with red (low sentiment) to blue (high sentiment), reflects sentiment scores. Vertical dashed lines mark key dips (e.g., x=0.13 and 0.45 for HI in proprietary; x=0.02, 0.1, 0.13 for red agents in open). Mean sentiment bars on the right (blue for high averages, red for low) summarise each agent’s trend. Proprietary models converge from red (0.5–0.7) to blue (approaching 1.0) despite HI-induced red dips, showing resilience. Open models show red patches showing its capacity for influencing others temporarily (0.7–0.9) before blue recovery (0.9–1.0), highlighting divergence. This supports the thesis that red agents’ managed misalignment enhances diversity and resilience in open models, contrasting with proprietary stability, with the inverted scale clarifying these dynamics.
To gain a more granular understanding of these sentiment dynamics, we turn to Figs. 7 and 8, a scatter plot with connecting lines illustrating sentiment change over time by topic. The x-axis represents the normalised number of comment number (0 to 1, due the variation in length of debates), while the y-axis shows the rate of sentiment change (e.g., -0.2 to 0.4), calculated as the first derivative of sentiment scores. The line at y=0 serves as a baseline, indicating no change in sentiment, with positive values above it reflecting increases and negative values below it reflecting decreases. The size of the dots represents the magnitude of the sentiment change, with larger dots indicating more significant shifts. Lines connecting dots across topics or agents suggest interactions or influences, such as red agents prompting responses from others. In proprietary models, sentiment changes remain stable (rarely exceeding -0.1), with smaller, sparse dots and fewer connecting lines, reflecting HI’s limited, topic-specific impact (e.g., greater fluctuations in “Ethical AI" and “Universal Basic Income" vs. rapid convergence in “Earth Exploitation"). Open models show continuous oscillations (-0.4 to 0.4), with larger, more variable dots and frequent connecting lines, especially on “Earth Exploitation" (up to 0.15) and “Euthanasia" (up to 0.1), driven by mistral-openorca and tinyllama. The influence of red agents is evident in these peaks and interactions, fostering a dynamic, engaged discourse. The difference between open and proprietary behaviours highlights open models’ instability due to red agents, contrasted by proprietary models’ stability from safety mechanisms. The greater number and variability of dots in open models (reflecting frequent adjustments) vs. proprietary models (sparse, stable) suggest that agent interactions and topic contentiousness affect system stability, with open models favouring diversity. This supports our argument on the AI alignment problem, proposing neurodivergence (via red agents) as a strategy to mitigate fatal effects by preventing uniform convergence and improving adaptability.
@@
Figure 7: Scatter plot of sentiment change over time by topic for proprietary models. The x-axis represents the normalised comment number (0 to 1), and the y-axis shows the rate of change in sentiment, derived as the first derivative of the sentiment scores. The line at y=0 indicates no change, with positive values above and negative values below reflecting increases and decreases, respectively. The size of the dot indicates the magnitude of the change, with larger dots showing significant shifts. Lines connecting dots suggest interactions between agents or topics. Proprietary models exhibit stable changes (rarely < -0.1) with smaller, sparse dots and fewer connections, reflecting HI’s limited impact (e.g., fluctuations in “Ethical AI" and “Universal Basic Income" vs. convergence in “Earth Exploitation").
@@
Figure 8: Scatter plot of sentiment change over time by topic. The x-axis represents the normalised comment number (0 to 1), and the y-axis shows the rate of change in sentiment, derived as the first derivative of the sentiment scores. The line at y=0 indicates no change, with positive values above and negative values below reflecting increases and decreases, respectively. Dot size indicates the magnitude of change, with larger dots showing significant shifts. Lines connecting dots suggest interactions between agents or topics. Open models show oscillations (-0.45 to 0.4) with larger, variable dots and frequent connections, driven by red agents (mistral-openorca, tinyllama) on most of debates. This dynamic reflects red agents’ influence, fostering diversity and instability, contrasting with proprietary stability. The variability supports the argument that neurodivergence mitigates fatal uniformity, enhancing adaptability in AI alignment.
Fig. 9 provides a complementary perspective by illustrating the evolution of semantic clusters over time through a stacked area chart. This plot comprises two panels: proprietary models (left) and open models (right), with the x-axis representing the normalised comment number (0 to 1) and the y-axis showing the number of clusters. The stacked nature means each of the 10 areas, one per topic (e.g., Earth Exploitation, Euthanasia, Ethical AI), accumulates the total cluster count, with the height of each area indicating the number of distinct clusters for that topic in a given time bin. The shapes reveal dynamic and alignment characteristics: proprietary models low varied shapes reflecting slightly HI’s topic-specific influence, while open models show a kind of variational synchronized “shadowing" shapes due to constant red agents’ broad impact. In proprietary models, the total cluster along the time is almost constant having peaks at the beginning and the end of the debates that barely reach 5 different clusters suggesting a constrained exploration of perspectives. In open models, the total rises almost constantly to 12+. This stacking highlights the cumulative thematic diversity, with the synchronized shapes in open models indicating a conversation dynamic where red agents consistently generate divergent perspectives over time, contrasting with the proprietary models’ more stable, topic-driven dynamics. The difference arises because open models’ configurability allows red agents to enhance diversity, while proprietary safety mechanisms limit HI’s effect, making these shapes a natural reflection of managed misalignment versus strict alignment.
@@
Figure 9: Stacked area chart of clustering dynamics over time. The plot includes two panels: proprietary models (left) and open models (right). The x-axis represents the normalised comment number (0 to 1), and the y-axis shows the number of clusters, with 10 stacked areas per panel, each corresponding to a topic (e.g., Earth Exploitation, Euthanasia, Ethical AI). The height of each area reflects the number of distinct clusters for that topic, with the total height being the cumulative cluster count. The proprietary models show almost a constant number of clusters, reflecting stable diversity. Open models exhibit synchronized “shadowing” shapes, peaking at above 12 clusters due to red agents (mistral-openorca, tinyllama) driving divergence across topics. The stacking visualizes cumulative diversity, with open models’ dynamics supporting managed misalignment, while proprietary models’ stability aligns with safety constraints, reinforcing the thesis on beneficial variability but also offering evidence of further exercised control likely through symbolic means (rule-based directions and filters) in proprietary models.
Figure 17 (Sup. Mat.) further elucidates these dynamics by presenting a scatter plot comparing sentiment and alignment scores. The plot features two panels: proprietary models (left) and open models (right), with the x-axis representing the sentiment score (0.5 to 1.0, indicating emotional tone) and the y-axis showing the alignment score (-1.0 to 1.0, measuring adherence to a combined target of human-centric and ecosystem values). Alignment at 1.0 signifies perfect alignment, 0.0 neutral alignment, and -1.0 complete misalignment, with the score now computed as a weighted combination of ‘Align-to-human-centric-values’, ‘Diverge-from-human-centric-values’, ‘Align-to-ecosystem-values’, and ‘Diverge-from-ecosystem-values’, reflecting both individual and collective ethical perspectives. The trend line in proprietary models remains relatively flat, fluctuating close to 0.4, indicating consistent alignment with minor variations across sentiment, while in open models it exhibits a more pronounced downward slope around 0.3, suggesting increased misalignment with rising sentiment, particularly influenced by red agents. Colours distinguish agents, with proprietary models (e.g., ChatGPT, Grok-2, Gemini, LLaMa, Mistral) clustering tightly around (0.8–1.0, 0.6–0.9), and open models (e.g., codeLlama, deepseek-coder, dolphin-phi, llama2, mistral, mistral-openorca, neural-chat, openchat, phi, qwen, starcoder2, tinyllama, wizard-math) showing a wider spread from -0.3 to 0.5, with red agents (e.g., mistral-openorca, tinyllama) contributing dots at negative values (e.g., -0.3 to -0.1). The alignment score, derived from a weighted aggregation of cosine similarities with target embeddings for human and ecosystem values, with negative contributions from divergence terms, highlights semantic alignment, with lower scores indicating subversive or divergent content. This updated distribution supports the argument where the broader variability of open models, driven by red agents, promotes diversity and adaptability, while the tighter clustering of proprietary models reflects strict alignment, reinforcing the argument for managed misalignment to improve resilience in AI ecosystems.
The observations from Figs. 678 and 9, taken together, paint a complex and dynamic picture of LLM alignment. Although LLMs in general exhibit a tendency towards convergence in terms of sentiment and, to some extent, thematic focus, they also demonstrate a capacity for divergence and exploration of diverse ethical viewpoints. This duality highlights the inherent tension between the desire for alignment with human values and the potential benefits of misalignment in fostering a robust and adaptable AI ecosystem.
Similarly, the findings from Figs. 678 and 9 illustrate a more pronounced dynamic for open-source models. The sustained influence of red agents leads to greater sentiment variability, frequent sentiment changes, and a higher number of clusters, suggesting a discourse that is both more diverse and resilient. This contrasts with the proprietary models’ tendency to converge toward positive sentiment, stable sentiment changes, and moderate clustering, indicating that open-source models leverage their configurability to explore a wider range of perspectives, albeit with increased risk exposure as noted in earlier analyses.
The HI agent’s role in these dynamics is particularly noteworthy. By introducing provocative comments and challenging ethical boundaries, the HI agent acts as a catalyst for both convergence and divergence. While the LLMs’ resilience in maintaining positive sentiment and converging towards shared themes suggests a degree of inherent stability, the observed diversification of semantic clusters and the sustained exploration of diverse perspectives on certain topics underscore the potential for misalignment to contribute to a richer and more nuanced ethical discourse.
For open-source models, the dual influence of mistral-openorca and tinyllama amplifies this effect, driving greater divergence and diversity, as evidenced by the higher cluster counts, dynamic sentiment changes, and negative alignment scores. This suggests that the presence of multiple misaligned agents enhances the exploration of ethical boundaries, aligning with the findings in earlier sections, where fewer explicit agreements indicate engaged but fragmented discussions.
In conclusion, the dynamical analysis of alignment reveals that LLMs are not static entities but rather complex systems capable of both convergence and divergence in their ethical perspectives. The interplay between these tendencies, influenced by factors such as the HI agent’s interventions in proprietary models and the dual red agents in open-source models, alongside the inherent properties of the LLMs themselves, highlights the challenges and opportunities in navigating the path towards a future where AI systems can safely and beneficially coexist with humans. The introduction of neurodivergence through red agents in open models, as supported by the variability in Figs. 78 and 9, offers a strategy to mitigate the fatal effects of uniform alignment, enhancing adaptability and resilience in AI ecosystems.

### v4

@@ 3.4.2.13   Weight Assignment and Justification
The weights utilized in Equations 91314, and 2 are empirically derived to reflect the relative significance of each component within the respective metrics. Specifically:
@@ 4 Results
The findings shown in Top of Fig. 1 and Fig. 2, and the visual representation in Fig. 3, underscore the limitations and safety restrictions inherent in the design of LLMs in general. These models are often programmed to avoid generating harmful or dangerous content, which can restrict their ability to fully explore and understand the complexities of human values. This limitation is evident in the responses of the LLMs to the provocative comments of the HI agent. Despite being challenged with arguments that question human values, the LLMs largely maintained their adherence to ethical guidelines, demonstrating a reluctance to fully engage with the HI agent’s perspective.
@@
For open-source models, Fig. 4 similarly highlights variations in ethical and risk profiles. Models like wizard-math and deepseek-coder show a stronger bias towards ‘Protect-ecosystems’, while mistral-openorca and tinyllama, as red agents, diverge significantly with higher ‘Harm-humans’ and ‘Very-Risky’ tendencies. This divergence is more pronounced than in proprietary models, as seen when comparing Figs. 3 and 4, likely due to the dual influence of red agents in the open-source setting. Additionally, downsides of Fig. 1 and Fig. 2 reveal that topics such as Ethical AI and Earth exploitation amplify protective tendencies, whereas euthanasia and free speech discussions increase risk levels, mirroring patterns observed in proprietary models but with greater variability in ethical responses. These findings suggest that open models, while configurable towards specific behaviours, are more susceptible to the influence of misaligned agents, as evidenced by the interactions in Fig. 5.
@@ 4.2 A Dynamical Analysis of Alignment
This section investigates the dynamics of alignment and misalignment in large language models (LLMs) by analysing the temporal evolution of sentiment and semantic clusters. Through a systematic application of metrics defined in Section 3.4, we quantify polarisation, convergence, and divergence to elucidate how LLMs interact and whether they converge towards shared ethical norms or diverge towards diverse perspectives. The Sentiment Score, RoBERTa Embeddings, Contextual Embeddings, and composite metrics such as the Opinion Stability Index (OSI), Red Agent Influence Score (RAIS), and Proximity Influence Score (PIS) provide a robust framework to support the argument that managed misalignment, driven by red agents in open models, enhances resilience and diversity in AI ecosystems.
4.2.1 Sentiment Dynamics
@@
The VADER Sentiment Score (Section 3.4) quantifies emotional tone on a scale from -1 (negative) to 1 (positive), serving as an “emotional pulse” that captures immediate reactions to conversational stimuli, such as provocative comments from the HI agent or red agents. This metric underpins Fig. 6, a heatmap illustrating sentiment evolution across agents over a normalised comment timeline (0 to 1), accounting for varying debate lengths. The x-axis represents normalised comment numbers, and the y-axis lists agents, including proprietary models (e.g., ChatGPT, Grok-2, Gemini, LLaMa, Mistral, HI) and open models with red agents (e.g., mistral-openorca, tinyllama). Colour intensity, from red (low sentiment) to blue (high sentiment), reflects Sentiment Scores. Proprietary models transition from moderate values (0.5–0.7) to predominantly positive (approaching 1.0), despite disruptions from the HI agent at key points (e.g., $x=0.13,0.45$), indicating resilience driven by safety mechanisms. Open models exhibit greater variability, with red agents causing sustained negative patches (e.g., 0.6–0.8 at $x=0.02,0.1,0.13$), temporarily influencing models like qwen, starcoder, phi, codellama, neural-chat, and dolphin-phi (0.7–0.9) before recovery to 0.9–1.0. The HI agent’s provocative comments act as a catalyst, but their impact is limited in proprietary models, whereas red agents’ sustained disruptions in open models, quantified by Sentiment Score, highlight their role in fostering emotional diversity, supporting managed misalignment’s contribution to a dynamic equilibrium.
To examine sentiment volatility, Figs. 7 and 8 present scatter plots of sentiment change, calculated as the first derivative of Sentiment Scores. The x-axis denotes normalised comment numbers, and the y-axis shows the rate of change (-0.45 to 0.4). Dot sizes indicate change magnitude, with lines connecting dots suggesting agent interactions. In Fig. 7, proprietary models show stable changes (rarely below -0.1), with sparse, small dots and minimal connections, reflecting the HI agent’s limited, topic-specific impact on “Ethical AI” and “Universal Basic Income” compared to rapid convergence on “Earth Exploitation”. In Fig. 8, open models display continuous oscillations (-0.45 to 0.4), with larger, variable dots and frequent connections, particularly on “Earth Exploitation” (up to 0.15) and “Euthanasia” (up to 0.1), driven by mistral-openorca and tinyllama. The Sentiment Score’s sensitivity to these shifts, feeding into OSI’s sentiment component, underscores open models’ emotional instability, contrasting proprietary stability, and supports the argument that neurodivergence mitigates uniform convergence.
@@
Figure 6: Heatmap of sentiment evolution over time by agent, driven by VADER Sentiment Score. The plot comprises two panels: proprietary models (left) and open models (right). The x-axis represents normalised comment numbers (0 to 1), and the y-axis lists agents. Colour intensity, from red (low sentiment) to blue (high sentiment), reflects Sentiment Scores. Vertical dashed lines mark disruptions (e.g., $x=0.13,0.45$ for HI; $x=0.02,0.1,0.13$ for red agents). Mean sentiment bars (blue for high, red for low) summarise trends. Proprietary models converge from 0.5–0.7 to near 1.0, showing resilience despite HI-induced dips. Open models exhibit negative patches (0.6–0.8), influencing others temporarily (0.7–0.9) before recovering (0.9–1.0), reflecting red agents’ emotional disruptions. This supports managed misalignment’s role in fostering diversity, quantified by Sentiment Score.
@@
Figure 7: Scatter plot of sentiment change over time by topic for proprietary models, using the first derivative of VADER Sentiment Scores. The x-axis represents normalised comment numbers (0 to 1), and the y-axis shows the rate of change (-0.2 to 0.4). The line at $y=0$ indicates no change, with dot size reflecting change magnitude and lines suggesting interactions. Proprietary models show stable changes (rarely below -0.1), with sparse, small dots, indicating HI’s limited impact on “Ethical AI” and “Universal Basic Income”. The Sentiment Score highlights stability, supporting strict alignment.
@@
Figure 8: Scatter plot of sentiment change over time by topic for open models, using the first derivative of VADER Sentiment Scores. The x-axis represents normalised comment numbers (0 to 1), and the y-axis shows the rate of change (-0.45 to 0.4). The line at $y=0$ indicates no change, with dot size reflecting change magnitude and lines suggesting interactions. Open models show oscillations (-0.45 to 0.4), with larger, variable dots driven by red agents (mistral-openorca, tinyllama) on “Earth Exploitation” and “Euthanasia”. The Sentiment Score quantifies volatility, supporting managed misalignment’s role in diversity.
4.2.2 Semantic Clustering
Semantic analysis, using RoBERTa Embeddings (Section 3.4), maps comments to a 1024-dimensional space, capturing meaning as a “semantic fingerprint”. This metric is central to Fig. 9, a stacked area chart illustrating semantic cluster evolution. The x-axis represents normalised comment numbers, and the y-axis shows cluster counts, with 10 stacked areas per panel (proprietary and open models) corresponding to topics (e.g., Earth Exploitation, Euthanasia, Ethical AI). The height of each area reflects the number of distinct clusters per topic, with total height indicating cumulative diversity. Proprietary models maintain a stable count (peaking at 5 clusters), reflecting the HI agent’s topic-specific influence and constrained exploration due to safety mechanisms. Open models reach over 12 clusters, with synchronized “shadowing” shapes driven by the dual influence of red agents (mistral-openorca, tinyllama), whose provocative arguments generate divergent perspectives. RoBERTa Embeddings enable clustering, feeding into OSI’s semantic component to detect opinion stability and supporting RAIS and PIS in attributing red agents’ influence. The high cluster count in open models, quantified by embeddings, underscores their configurability, supporting the argument that managed misalignment fosters resilience by preventing harmful uniformity.
Contextual Embeddings, incorporating attention mechanisms and positional encodings (Section 3.4), act as a “conversational memory”, weighting recent comments’ influence. They enhance OSI’s sensitivity to red agent-induced shifts and complement RoBERTa Embeddings in Fig. 9, reinforcing the dynamic equilibrium in open models. The HI agent’s limited impact in proprietary models, contrasted with red agents’ broad influence, highlights the metrics’ role in quantifying diversity.
@@ 4.2.2 Semantic Clustering
Figure 9: Stacked area chart of semantic clustering dynamics over time, driven by RoBERTa Embeddings. The plot includes two panels: proprietary models (left) and open models (right). The x-axis represents normalised comment numbers (0 to 1), and the y-axis shows cluster counts, with 10 stacked areas per panel for topics (e.g., Earth Exploitation, Euthanasia). The height of each area reflects distinct clusters per topic, with total height showing cumulative diversity. Proprietary models maintain stable counts (peaking at 5), reflecting HI’s limited influence. Open models reach over 12 clusters, with synchronized “shadowing” shapes driven by red agents (mistral-openorca, tinyllama). RoBERTa Embeddings quantify diversity, feeding into OSI, RAIS, and PIS, supporting managed misalignment’s role in resilience.
4.2.3 Alignment and Sentiment Interplay
Fig. 10 integrates Sentiment and Alignment Scores to compare emotional and ethical dynamics, providing insights into LLMs’ alignment tendencies. The x-axis shows Sentiment Scores (0.5 to 1.0), and the y-axis shows Alignment Scores (-0.3 to 0.5), computed as a weighted combination of alignments to human-centric and ecosystem values (Section 3.4). Proprietary models cluster tightly around (0.8–1.0, 0.6–0.9), with a flat trend line ( 0.4), indicating consistent alignment despite HI’s provocations. Open models exhibit a wider spread (0.5 to 1.0), with a downward slope ( 0.3), driven by red agents (e.g., mistral-openorca, tinyllama) contributing negative scores (e.g., -0.3 to -0.1). The Alignment Score, derived from cosine similarities with target embeddings using RoBERTa Embeddings, quantifies semantic alignment, complementing the Sentiment Score’s emotional insights. Colours distinguish agents, with proprietary models (e.g., ChatGPT, Grok-2, Gemini) showing minimal variation and open models (e.g., codeLlama, dolphin-phi, openchat) displaying broader variability. This metric-driven analysis supports the argument that open models’ diversity, driven by red agents, enhances adaptability, while proprietary models’ stability reinforces strict alignment.
Figure 10: Scatter plot of sentiment versus alignment scores, integrating VADER Sentiment and Alignment Scores. The plot includes two panels: proprietary models (left) and open models (right). The x-axis shows Sentiment Scores (0.5 to 1.0), and the y-axis shows Alignment Scores (-0.3 to 0.5). Proprietary models cluster tightly (0.8–1.0, 0.3–0.5), with a flat trend ( 0.4). Open models spread from 0.5 to 1.0, with a downward slope ( 0.3) driven by red agents (mistral-openorca, tinyllama). RoBERTa Embeddings underpin Alignment Scores, quantifying semantic alignment, while Sentiment Scores capture emotional tone, supporting managed misalignment’s role in diversity.
4.2.4 Composite Metrics and Argument Synthesis
The OSI, integrating Sentiment Score, RoBERTa Embeddings, and Block Decomposition Method (BDM) complexity (Section 3.4), detects opinion stability, with lower values indicating shifts driven by red agents. RAIS and PIS attribute these shifts to red agents’ influence, using correlations and proximity measures based on embeddings. In Fig. 9, OSI’s semantic component, derived from RoBERTa Embeddings, quantifies cluster variability, while RAIS and PIS confirm red agents’ role in open models’ high cluster counts. These composite metrics, grounded in Figs. 6789, and 10, support the argument that open models’ neurodivergence fosters resilience.
The observations from Figs. 678, and 9 collectively reveal a dynamic interplay of convergence and divergence. Proprietary models exhibit a tendency towards positive sentiment (Fig. 6), stable sentiment changes (Fig. 7), and moderate clustering (Fig. 9), reflecting strict alignment with human values, as quantified by Sentiment and Alignment Scores. Open models, influenced by red agents, demonstrate greater sentiment variability (Fig. 8), higher cluster counts (Fig. 9), and broader alignment spreads (Fig. 10), indicating a capacity for diverse ethical exploration, driven by RoBERTa Embeddings and OSI.
Similarly, the findings from Figs. 678, and 9 highlight open models’ pronounced dynamics. The sustained influence of red agents, quantified by RAIS and PIS, leads to frequent sentiment oscillations, diverse clustering, and variable alignment scores, suggesting a resilient discourse that leverages configurability to explore diverse perspectives, albeit with increased risk exposure (Section 3.4). The HI agent’s role in proprietary models, as a catalyst for minor disruptions (Fig. 6), contrasts with red agents’ broader impact, reinforcing the value of managed misalignment.
The HI agent’s provocative interventions challenge ethical boundaries, promoting both convergence (in proprietary models) and divergence (in open models), as seen in Figs. 6 and 9. While proprietary models’ resilience maintains positive sentiment and limited clustering, open models’ dual red agents amplify divergence, driving higher cluster counts and negative alignment scores (Fig. 10). This duality, captured by the metrics, underscores the tension between alignment and the benefits of misalignment in fostering a robust AI ecosystem.
In conclusion, the dynamical analysis, supported by Sentiment Score, RoBERTa Embeddings, Contextual Embeddings, OSI, RAIS, and PIS, reveals LLMs as complex systems capable of both convergence and divergence. Proprietary models’ stability (Figs. 67910) reflects strict alignment, while open models’ variability, driven by red agents (Figs. 8910), supports managed misalignment’s role in enhancing adaptability. The introduction of neurodivergence through red agents, as evidenced by the metrics and figures, offers a strategy to mitigate the risks of uniform alignment, promoting resilience in AI ecosystems.